### v1

# Deep Learning-Based CSI Feedback for Wi-Fi Systems With Temporal Correlation

Abstract
To achieve higher throughput in next-generation Wi-Fi systems, a station (STA) needs to efficiently compress channel state information (CSI) and feed it back to an access point (AP). In this paper, we propose a novel deep learning (DL)-based CSI feedback framework tailored for next-generation Wi-Fi systems. Our framework incorporates a pair of encoder and decoder neural networks to compress and reconstruct the angle parameters of the CSI. To enable an efficient finite-bit representation of the encoder output, we introduce a trainable vector quantization module, which is integrated after the encoder network and jointly trained with both the encoder and decoder networks in an end-to-end manner. Additionally, we further enhance our framework by leveraging the temporal correlation of the angle parameters. Specifically, we propose an angle-difference feedback strategy which transmits the difference between the current and previous angle parameters when the difference is sufficiently small. This strategy accounts for the periodicity of the angle parameters through proper preprocessing and mitigates error propagation effects using novel feedback methods. We also introduce a DL-based CSI refinement module for the AP, which improves the reconstruction accuracy of the angle parameters by simultaneously utilizing both the previous and current feedback information. Simulation results demonstrate that our framework outperforms the standard method employed in current Wi-Fi systems. Our results also demonstrate significant performance gains achieved by the angle-difference feedback strategy and the CSI refinement module.
Index Terms:
Wi-Fi, channel state information (CSI) feedback, deep-learning-based feedback, angle-difference feedback, temporal correlation
I Introduction
Recently, wireless networks have dramatically advanced to meet the growing demand for higher data rates. Specifically, the IEEE 802.11 task group has been developing a novel extremely high throughput (EHT) amendment for wireless local area networks (WLANs) [1]. With developing an advanced multiple-input multiple-output (MIMO) techniques, it aims to service much higher throughput and reliable communications. To achieve higher throughput in next-generation Wi-Fi systems utilizing MIMO techniques, acquiring an accurate downlink channel state information (CSI) is indispensable for the access point (AP) to transmit data to the station (STA). The typical CSI acquisition process in the system is initiated by the AP’s transmitting the null data packet announcement (NDPA) frame to STA. NDPA serves the role of notifying the STA of all necessary probing details of MIMO transmission. Then, a null data packet (NDP) is transmitted to the STA for estimating the downlink CSI. After the channel estimation, the STA compresses and feeds the CSI back to the AP, in the form of the compressed beamforming report (CBR). The compression is performed by extracting angle parameters after performing singular value decomposition (SVD) of the MIMO channel. The aforementioned protocol in the Wi-Fi systems is summarized in Fig. 1. Another approach to acquire the downlink CSI is to leverage channel reciprocity. In this process, the downlink CSI is inferred from the uplink CSI with additional calibrations. However, this approach is excluded from IEEE 802.11ac standard [2], indicating its verified limitation.
Figure 1: An illustration of a communication protocol in Wi-Fi systems.
The CSI feedback methodologies have been an important problem in cellular networks as well. Among various methodologies, deep learning (DL)-based CSI feedback methods have been proposed for the cellular networks in recent years [3, 4, 5, 6, 7, 8]. By properly preprocessing the CSI and carefully designing neural networks, these methods have demonstrated their effectiveness in reducing CSI feedback overhead, outperforming traditional non-DL-based approaches like compressed sensing and static codebook-based feedback[9, 10]. A typical approach of the DL-based CSI feedback is to compress a preprocessed CSI into a low-dimensional latent vector by using an encoder network. Then, the CSI is reconstructed by using a decoder network from a received latent vector. To enable the transmission of the latent vector with a finite number of bits, one possible approach is to apply scalar quantization (SQ) to each latent element. For example, non-uniform SQ methods were explored in [11, 12] to represent the latent vector with a limited number of bits for CSI feedback in cellular networks. However, the application of SQ to DL-based CSI feedback has inherent limitations. First, SQ disregards the correlations between latent elements, resulting in suboptimal performance when these elements are correlated [13]. Second, SQ cannot allocate fewer than one quantization bit per latent entry, which imposes constraints on the dimensionality of the latent vector for a given feedback overhead and ultimately limits the feedback efficiency. These limitations can be resolved by applying a vector quantization (VQ) in the latent space. The application of VQ to the latent representation allows for modeling the correlation among latent entries. Furthermore, the joint optimization of the VQ codebook with the entire model facilitates quantization that effectively reflects the distribution of the latent vectors, thereby enhancing the efficiency of the learned representations [14]. Motivated by this, DL-based CSI feedback methods incorporating VQ have been proposed for cellular networks in [15, 16, 17]. In [15], it was reported that the DL-based CSI feedback with the VQ framework can achieve a better CSI reconstruction performance than that of DL-based CSI feedback with the SQ framework in cellular networks.
To further improve the efficiency of DL-based CSI feedback, some studies have leveraged the temporal correlation of CSI, which naturally arises in practical communication scenarios involving devices with low to moderate mobility. A common approach in this direction is to employ neural network architectures specifically designed for time-series inputs. For instance, it was reported in [18, 19] that incorporating long short-term memory (LSTM) [20] enables the model to capture temporal correlation, enhancing feedback performance. On the other hand, an alternative approach to leveraging the temporal correlation of CSI is to exploit the inherent sparsity in the differences between sequential CSI instances. This approach stems from the intuition that sequential instances are likely to exhibit similar values, resulting in sparse differences which is more efficient form to compress and reconstruct. The sparsity in CSI differences was leveraged in [22, 21], demonstrating an efficacy of this approach in improving the performance of the CSI feedback in cellular networks.
Unlike DL-based CSI feedback in cellular networks, limited attention has been given to DL-based CSI feedback for Wi-Fi systems [23, 24]. In [23], the angle parameters of downlink CSI were separately compressed using hyperparameterized preprocessing and two distinct fully connected networks (FCNs). In [24], a convolutional neural network (CNN)-based model with a channel attention module was used to compress a right singular matrix in SVD, rather than the angle parameters. However, these studies employ simple quantization techniques without jointly training the quantization module with the encoder and decoder networks. Moreover, none of the existing research has leveraged the temporal correlation of CSIs in DL-based CSI feedback specifically for Wi-Fi systems. The potential benefits of the joint training of the quantization module and the utilization of temporal correlation in CSIs remain unexplored, despite their critical importance in realizing the full potential of DL-based CSI feedback in next-generation Wi-Fi systems.
To fill this research gap, in this paper, we propose novel DL-based CSI feedback and refinement frameworks that pioneer the use of a trainable VQ module and the exploitation of temporal correlation in CSIs for Wi-Fi systems. The major contributions in this paper are summarized as follows:
  • We introduce an initial feedback strategy for our framework, which integrates a trainable VQ module with the DL-based CSI feedback process in Wi-Fi systems. Our strategy basically facilitates joint training of the VQ module with encoder and decoder networks. This joint training enables an efficient finite-bit representation of the latent vector, improving feedback efficiency. Additionally, we determine an appropriate input form for the encoder network by aggregating angle parameters specified in the IEEE 802.11 standard.
  • We propose a novel angle-difference feedback strategy for Wi-Fi systems that leverages the temporal correlation of CSIs, an aspect that has never been explored in DL-based CSI feedback for Wi-Fi systems. The core idea of our framework is to feed back the difference between the current and previous angle parameters when the difference is sufficiently small. To ensure an appropriate input for the angle-difference encoder network, we introduce a preprocessing step designed to capture the periodicity of angle parameters in Wi-Fi systems. Moreover, to address the challenge of error propagation inherent in the angle-difference feedback, we design two VQ architectures, referred to as parallel VQ and unified VQ, each tailored to compensate for residual errors from the previous feedback.
  • We propose a novel CSI refinement framework for Wi-Fi systems that enhances reconstruction performance at the AP by leveraging the temporal correlation of CSIs. A key ingredient of this framework is a CSI refinement module designed to refine the angle parameters at each time instance using both the previous and current feedback information. To further improve its performance, we also introduce a recursive refinement strategy, where the refined angle parameter is utilized as an input for subsequent CSI refinement processes. These advancements provide a new approach to exploiting temporal correlations in CSIs on the AP side as well, further improving the efficiency and accuracy of CSI reconstruction in Wi-Fi systems.
  • Using extensive simulations, we validate the effectiveness of the proposed CSI feedback and refinement frameworks in Wi-Fi systems with temporally-correlated channels. Our evaluations leverage a CSI dataset generated from the IEEE 802.11be standard configuration [1]. The simulation results reveal that the proposed CSI feedback, angle-difference feedback, and CSI refinement frameworks consistently achieve significant performance gains across diverse scenarios. Furthermore, we demonstrate that the proposed frameworks outperforms the standard feedback method employed in current Wi-Fi systems, both in terms of CSI reconstruction accuracy and net throughput performance. Through these results, we demonstrate the potential of our frameworks in advancing the efficiency of next-generation Wi-Fi systems.
This work builds upon our previous study [25], where we introduced only the DL-based angle-difference CSI feedback framework. In the current work, we newly propose a DL-based CSI refinement strategy to further improve the feedback efficiency by leveraging the temporal correlation of CSIs on the AP side. To refine the angle parameters effectively at each time step, we design a CSI refinement module and a recursive strategy that utilize both previous and current feedback. With this extension, we enhance the simulation study to demonstrate the superiority of our framework in Wi-Fi systems. Our simulation results demonstrate the superiority of the proposed CSI refinement framework in enhancing the reconstruction performance and efficiency of the CSI feedback in Wi-Fi systems.
[…]
II System Model and preliminary
II-A Downlink MIMO Wi-Fi Systems
Consider a downlink Wi-Fi system in which an AP equipped with $N_{t}$ transmit antennas transmits data to a STA equipped with $N_{r}$ receive antennas. The system employs orthogonal frequency division multiplexing (OFDM) with $N_{c}$ subcarriers. To facilitate channel estimation at the STA, the AP transmits the NDP to the STA and a received NDP can be expressed as
$\mathbf{y}[k]=\mathbf{H}[k]\mathbf{s}[k]+\mathbf{n}[k],\quad k=0,1,...,N_{c}-1,$ (1)
where $\mathbf{y}[k]\in\mathbb{C}^{N_{r}},\mathbf{H}[k]\in\mathbb{C}^{N_{r}\times N_{t}},\mathbf{s}[k]\in\mathbb{C}^{N_{t}},$ and $\mathbf{n}[k]\in\mathbb{C}^{N_{r}}$ represent the received NDP, channel matrix, transmitted NDP, and additive noise vector respectively. Then, the STA estimates the CSI (i.e., channel frequency responses (CFRs) associated with all subcarriers) using the received NDP. The channel estimation is assumed to be perfect in this work. In the IEEE 802.11 standard protocols, to extract beamforming matrices, the STA computes the SVD of each CFR as
$\mathbf{H}[k]=\mathbf{U}[k]\mathbf{\Sigma}[k]\tilde{\mathbf{V}}^{\text{H}}[k],$ (2)
where $\mathbf{U}[k]\in\mathbb{C}^{N_{r}\times N_{r}}$ and $\mathbf{\tilde{V}}[k]\in\mathbb{C}^{N_{t}\times N_{t}}$ are unitary matrices, and $\mathbf{\Sigma}[k]\in\mathbb{C}^{N_{r}\times N_{t}}$ is a diagonal matrix. In the SVD beamforming, $\mathbf{U}[k]$ and $\mathbf{\tilde{V}}[k]$ are utilized to design receive and transmit beamforming matrix, respectively. To this end, in the IEEE 802.11 protocols, the first $N_{s}$ columns of $\tilde{\mathbf{V}}[k]$, denoted as $\overline{\mathbf{V}}[k]\in\mathbb{C}^{N_{t}\times N_{s}}$, are compressed and fed back to the AP, where $N_{s}$ represents the number of data streams.
Now, we present the CSI compression method employed in the IEEE 802.11 standard, which is based on the Givens rotation with angle parameters $(\phi,\psi)$[27, 28, 29]. First, the phase offset of the last row of $\overline{\mathbf{V}}[k]$ is eliminated as
$\mathbf{V}[k]=\mathbf{\overline{V}}[k]\mathbf{\tilde{D}}^{*}[k],$ (3)
where $\mathbf{\tilde{D}}[k]\in\mathbb{C}^{N_{t}\times N_{t}}$ is a diagonal matrix, whose diagonal components are composed of phase coefficients of the last row of $\mathbf{\overline{V}}[k]$. Note that it is sufficient for the AP to use the transmit beamforming with $\mathbf{V}[k]$, since $\mathbf{\tilde{D}}[k]$ is a diagonal matrix and exchangable with $\mathbf{\Sigma}[k]$, so that it can be equalized by STA. This ensures that last row of $\mathbf{V}[k]$ consists of non-negative real values.
Next, $\mathbf{V}[k]$ is rewritten as the product of the Givens rotation matrix as
$\mathbf{V}[k]$
$=\left(\prod^{{\rm min}(N_{s},N_{t}-1)}_{i=1}\left(\mathbf{D}_{i}\big{(}\bm{\phi}_{i}^{(k)}\big{)}\prod^{N_{t}}_{l=i+1}\mathbf{G}_{li}^{\sf T}\big{(}\psi_{li}^{(k)}\big{)}\right)\right)\mathbf{I}_{N_{t}\times N_{s}},$ (4)
where $\mathbf{D}_{i}\left(\bm{\phi}_{i}\right)$ is an $N_{t}\times N_{t}$ diagonal matrix defined as
$\mathbf{D}_{i}(\bm{\phi}_{i})=\begin{bmatrix}\mathbf{I}_{i-1}&0&0&\cdots&0\\ 0&e^{j\phi_{i,i}}&0&\cdots&0\\ 0&0&\ddots&0&0\\ 0&0&\cdots&e^{j\phi_{N_{t}-1,i}}&0\\ 0&0&\cdots&0&1\end{bmatrix},$ (5)
and $\mathbf{G}_{li}\left(\psi_{li}\right)$ is an $N_{t}\times N_{t}$ Givens rotation matrix defined as
$\mathbf{G}_{li}(\psi_{li})=\begin{bmatrix}\mathbf{I}_{i-1}&0&0&0&0\\ 0&\text{cos}\psi_{li}&0&\text{sin}\psi_{li}&0\\ 0&0&\mathbf{I}_{l-i-1}&0&0\\ 0&-\text{sin}\psi_{li}&0&\text{cos}\psi_{li}&0\\ 0&0&0&0&\mathbf{I}_{N_{t}-l}\end{bmatrix}.$ (6)
Also, $\mathbf{I}_{m}$ represents an $m\times m$ identity matrix and $\mathbf{I}_{m\times n}$ represents an $m\times n$ ($m\geq n$) matrix which corresponds to zero-padded version of $\mathbf{I}_{m}$. In the case of $N_{r}=N_{t}=N_{s}=3$, for instance, $\mathbf{V}[k]$ can be decomposed as
$\mathbf{V}[k]=$ $\begin{bmatrix}e^{j\phi^{(k)}_{11}}&0&0\\ 0&e^{j\phi^{(k)}_{21}}&0\\ 0&0&1\end{bmatrix}\times\begin{bmatrix}\text{cos}\psi^{(k)}_{21}&\text{sin}\psi^{(k)}_{21}&0\\ -\text{sin}\psi^{(k)}_{21}&\text{cos}\psi^{(k)}_{21}&0\\ 0&0&1\end{bmatrix}^{\rm T}$
×$\times\begin{bmatrix}\text{cos}\psi^{(k)}_{31}&0&\text{sin}\psi^{(k)}_{31}\\ 0&1&0\\ -\text{sin}\psi^{(k)}_{31}&0&\text{cos}\psi^{(k)}_{31}\end{bmatrix}^{\rm T}\times\begin{bmatrix}1&0&0\\ 0&e^{j\phi^{(k)}_{22}}&0\\ 0&0&1\end{bmatrix}$
×$\times\begin{bmatrix}1&0&0\\ 0&\text{cos}\psi^{(k)}_{32}&\text{sin}\psi^{(k)}_{32}\\ 0&-\text{sin}\psi^{(k)}_{32}&\text{cos}\psi^{(k)}_{32}\\ \end{bmatrix}^{\rm T}\times\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix},$ (7)
with angle parameters $[\phi^{(k)}_{11},\phi^{(k)}_{21},\psi^{(k)}_{21},\psi^{(k)}_{31},\phi^{(k)}_{22},\psi^{(k)}_{32}]^{\rm T}$ extracted for the $k$-th subcarrier. From (II-A), the angle parameters $\{\phi_{j,i}\}$ and $\{\psi_{l,i}\}$ can be extracted using Algorithm 1.
1. Parameters: $\bm{\phi}^{(k)}_{i}=[\phi^{(k)}_{i,i},...,\phi^{(k)}_{N_{t}-1,i}]$, $\psi^{(k)}_{li}$
2. Extraction:
for $k=0,...,N_{c}-1$ do
       $[\mathbf{U},\mathbf{\Sigma},\mathbf{\tilde{V}}]\leftarrow\text{SVD}(\mathbf{H}[k])$;
       $\mathbf{\overline{V}}\leftarrow[\mathbf{\tilde{V}}]_{:,1:N_{s}}$;
       $\mathbf{\tilde{D}}\leftarrow{\rm diag}({\rm exp}(j\times{\rm phase}([\mathbf{\overline{V}}]_{N_{t},:})))$;
       $\mathbf{V}\leftarrow\mathbf{\overline{V}}\mathbf{\tilde{D}}^{*}$;
       for $i=1,...,{\rm min}(N_{s},N_{t}-1)$ do
             $\bm{\phi}^{(k)}_{i}\leftarrow{\rm phase}([\mathbf{V}]_{i:N_{t}-1,i})$;
             $\mathbf{V}\leftarrow\mathbf{D}_{i}(\bm{\phi}^{(k)}_{i})^{*}\mathbf{V}$;
             for $l=i+1,...,N_{t}$ do
                   $\psi^{(k)}_{li}={\rm tan}^{-1}\left(\frac{{\rm Re}\{[\mathbf{V}]_{l,i}\}}{{\rm Re}\{[\mathbf{V}]_{i,i}\}}\right)$;
                   $\mathbf{V}\leftarrow\mathbf{G}_{li}(\psi^{(k)}_{li})^{*}\mathbf{V}$;
             end for
       end for
end for
Algorithm 1 Algorithm for extracting $\phi$ and $\psi$ from $\mathbf{\tilde{V}}[k]$
Last, the angle parameters are uniformly quantized using $b_{\phi}$ and $b_{\psi}$ bits, respectively. The quantized parameters are expressed as
$\hat{\phi}_{j,i}$$\in\left\{\frac{n\pi}{2^{b_{\phi}-1}}+\frac{\pi}{2^{b_{\phi}}}\ \big{|}\ n=1,2,...,2^{b_{\phi}-1}\right\},$ (8)
$\hat{\psi}_{l,i}$$\in\left\{\frac{n\pi}{2^{b_{\psi}+1}}+\frac{\pi}{2^{b_{\psi}+2}}\ \big{|}\ n=1,2,...,2^{b_{\psi}-1}\right\}.$ (9)
These parameters are fed back to the AP in order of $[...,\ \hat{\phi}^{(k)}_{i:N_{t}-1,i},\ \hat{\psi}^{(k)}_{i+1:N_{t},i},\ ...]^{\rm T}$ for the $k$-th subcarrier. The aforementioned feedback method provides satisfactory performance when the number of feedback bits is sufficient, which is why it has been adopted in the IEEE 802.11 standard to date. However, when feedback overhead is limited, this method suffers from significant performance degradation, potentially creating a bottleneck in next-generation Wi-Fi systems. This limitation naturally motivates the development of a new CSI feedback method based on DL.
II-B DL-based CSI Feedback Process
To implement a typical DL-based CSI feedback method for Wi-Fi systems, the STA employs an encoder network to compress the CSI input into the form of a latent vector with dimension $M$. Meanwhile, the AP employs a decoder network to reconstruct the original CSI from the STA’s feedback.
The STA initiates a CSI feedback process by utilizing the CSI input $\mathbf{X}$ for the encoder network. Note that our choice for this input will be discussed in Sec. III-A. This yields the latent vector ${\bf z}$ expressed as ${\bf z}=f_{\text{enc}}(\mathbf{X})$, where $f_{\text{enc}}$ represents the encoder network. Typically, the latent vector ${\bf z}$ is considered as a STA’s feedback that needs to be transmitted to the AP. Upon the reception of the feedback, the AP reconstructs the CSI by utilizing the latent vector as an input of the decoder network $f_{\text{dec}}$. In practice, the latent vector ${\bf z}$ is transformed into a finite-length bit sequence before being transmitted to the AP. This transformation is achieved through quantizing the latent vector $\mathbf{z}$ under the constraint on feedback overhead. Subsequently, the AP reconstructs the CSI by utilizing the quantized latent vector, denoted as ${\bf z}_{q}$, as an input for the decoder network, i.e., $\mathbf{\hat{X}}=f_{\text{dec}}(\mathbf{z}_{q})$.
III Proposed DL-based CSI Feedback Framework
In this section, we propose a novel DL-based CSI feedback framework which utilizes VQ approach to realize the efficient finite-bit feedback of angle parameters in Wi-Fi systems, while leveraging the temporal correlation of these parameters. First, we introduce an initial CSI feedback method that employs VQ to efficiently compress and feedback individually obtained CSI. Then, we present an angle-difference feedback method that leverages temporal correlation among sequentially acquired CSI, in conjunction with the proposed VQ structures. Specifically, we propose two types of VQ structures—parallel VQ and unified VQ—that not only exploit temporal correlation but also address the issue of compensating for the quantization error from the previous feedback.
III-A Initial Feedback Method: DL-based CSI Feedback with VQ
Building on the effectiveness of VQ in facilitating finite-bit representations of latent vectors, we adopt it as the foundation of the inital CSI feedback framework. Our framework utilizes a trainable quantization codebook placed in the latent space and jointly trains the encoder, codebook, and decoder using a loss function that captures both the quantization and reconstruction errors. To avoid a significant computational complexity required for VQ, we adopt the product VQ structure in [26]. Specifically, we first divide the latent vector ${\bf z}$ into $N$ sub-vectors each with dimension $D$ ($M=N\times D$) and then use a $D$-dimensional codebook for quantizing each sub-vector separately, as illustrated in Fig. 2. Let $\mathcal{B}$ be a vector codebook using $B$ bits which consists of $2^{B}$ $D$-dimensional codewords, namely $\{\mathbf{b}_{k}\}^{2^{B}}_{k=1}$. Also, let ${\bf z}_{i}$ be the $i$-th sub-vector of ${\bf z}$, defined as ${\bf z}_{i}=[z_{(i-1)D+1},\cdots,z_{iD}]$, where $z_{j}$ is the $j$-th entry of ${\bf z}$. Then each sub-vector ${\bf z}_{i}$ is quantized to $\mathbf{z}_{q,i}$ using the codebook $\mathcal{B}$ according to the minimum Euclidean distance criterion, i.e.,
$\mathbf{z}_{q,i}=\operatornamewithlimits{argmin}_{\mathbf{b}_{k}\in\mathcal{B}}\|\mathbf{z}_{i}-\mathbf{b}_{k}\|.$ (10)
For jointly training the encoder, codebook, and decoder, we adopt the loss function in [14], given by
$\mathcal{L}_{\text{vq}}=d(\mathbf{X},\mathbf{\hat{X}})+\left\|\text{sg}(\mathbf{z})-\mathbf{z}_{q}\right\|^{2}+\beta\left\|\mathbf{z}-\text{sg}(\mathbf{z}_{q})\right\|^{2},$ (11)
where $d(\cdot,\cdot)$ is a distortion measure which calculates the difference between two inputs with a certain criterion. Also, sg($\cdot$) is the ‘stop-gradient’ operator, which treats an input as a constant, allowing it to be ignored in gradient descent procedures with a given loss structure. In particular, the last two terms in (11) depend on a quantization loss, aiming to minimize the difference between $\mathbf{z}$ and $\mathbf{z}_{q}$. The third term in (11) is called a commitment loss which is regularized by a hyperparameter $\beta$ [14]. After calculating the quantization loss, a gradient correction of the decoder input is performed as $\mathbf{z}_{q}\leftarrow\mathbf{z}+\text{sg}(\mathbf{z}_{q}-\mathbf{z})$.
Figure 2: An illustration of our DL-based initial CSI feedback with VQ.
In our method, we choose an angle parameter $\mathbf{\Phi}\in\mathbb{R}^{2\times N_{a}\times N_{c}}$ as the input $\mathbf{X}$ of the encoder network, which consists of the angle parameters $\{\phi_{i,j}\}$ and $\{\psi_{l,i}\}$, where $N_{a}={\rm max}(\sum^{N_{t}-1}_{i={\rm max}(N_{t}-N_{s},1)}i,1)$ represents the number of each angle parameter per subcarrier. We aggregate the angle parameters in $\mathbf{\Phi}$ in the same order as specified in the IEEE 802.11 standard. The rationale behind this input choice is that we can exploit the abstracted information in the beamforming matrix with orthogonal columns, while effectively accounting for the correlations between these elements. Otherwise, if we were to use ${\bf V}$ as the input instead, the input dimension would typically be larger, and an additional enforcement process would be required to ensure the output has orthogonal columns explicitly. This rationale is numerically justified in Sec. V-B.
Remark (Comparison with Existing DL-based CSI Feedback): Our approach is differentiated from existing DL-based CSI feedback methods for Wi-Fi systems, which are those in [23, 24], with several viewpoints. First, we employ a trainable VQ module which facilitate an efficient finite-bit representation of the latent vector compared to the quantization techniques in [23, 24]. We also consider the angle parameters $\{\phi_{i,j}\}$ and $\{\psi_{l,i}\}$ as an input for the encoder network, resulting in a more efficient input representation than the inputs considered in the existing methods.
Figure 3: An illustration of preprocessing on $\mathbf{\Phi}^{\rm diff}_{t+1}$ designed to capture the periodicity of angle parameters.
III-B Core Idea: Angle-Difference Feedback
In practical systems, the STA connected to the AP periodically feeds the CSI back to the AP. In this scenario, the CSIs continuously evolve over time and therefore have temporal correlations up to a certain level. Inspired by this fact, in this subsection, we present a core idea of angle-difference feedback strategy which further enhances the feedback efficiency by exploiting the temporal correlation of the CSIs.
A simple yet effective approach to exploit the temporal correlation of the CSIs is to feed back the difference between the current and previous angle parameters to the AP. If these angle parameters are temporally correlated, their difference would be sparse. This sparsity makes it more efficient to compress the sparse matrix and then sum it with the original matrix after reconstruction. To incorporate this idea into our CSI feedback framework in Sec. III-A, we consider two feedback phases: (i) an initial CSI feedback phase, and (ii) an angle-difference feedback phase. In the initial feedback phase, the angle parameter $\mathbf{\Phi}_{t}$ at time $t$ is compressed using an initial encoder, referred to as the Type-I encoder $f_{\rm enc1}$. The corresponding decoder, referred to as the Type-I decoder $f_{\rm dec1}$, is then utilized at the AP to reconstruct the initial angle parameter. In the angle-difference feedback phase, the angle difference at time $t$, defined as $\mathbf{\Phi}^{\rm diff}_{t}=\mathbf{\Phi}_{t}-\mathbf{\Phi}_{t-1}$, is compressed using an angle-difference encoder, namely Type-II encoder $f_{\rm enc2}$. The corresponding decoder, referred to as the Type-II decoder $f_{\rm dec2}$, is used to reconstruct the angle difference $\mathbf{\Phi}^{\rm diff}_{t}$. During this phase, the angle parameter is reconstructed by adding the reconstructed angle difference to the previous angle parameter.
III-C Preprocessing for Angle-Difference Feedback
Using the angle difference $\mathbf{\Phi}^{\rm diff}_{t}$ directly as the input to the Type-II encoder $f_{\rm enc2}$ fails to account for the periodicity of angle parameters in Wi-Fi systems. To enable the model to effectively capture actual variations in these angle parameters, in this subsection, we design a proper preprocessing to capture the periodicity of the angle parameters in Wi-Fi systems. To achieve this goal, we define a preprocessing function $h(x)$, applied to each element $[\mathbf{\Phi}^{\rm diff}_{t}]_{m,n}\ \forall m,n$, as